a box has a square bottom and is 6cm tall. the total surface area is 378 cm^2. find the volume of the box.
Answer by richwmiller(17219)

2x^2+24x-378=0
2 (x^2+12 x-189) = 0
2 (x-9) (x+21) = 0
x=9
9*9*6=486 cubic cm volume
check
9*9*2=162
4*9*6=24*9=216
162+216=378
ok